OpenStudy (anonymous):

Assume a jar has 7 red marbles and 5 black marbles. Draw out 3 marbles with and without replacement. Find the requested probabilities. (c) P(one red and two black marbles) With replacement , without replacement . (d) P(red on the first draw and black on the second draw and black on the third draw) With replacement , without replacement .

OpenStudy (anonymous):

What does with replacement mean?

OpenStudy (anonymous):

putting back in the jar

OpenStudy (anonymous):

Oh. I dunno about that, but without replacement is pretty easy.

OpenStudy (anonymous):

i understand more of the with replacement but if you can help with the without replacement that would be good

OpenStudy (anonymous):

For the first question of one red and two black we just want to know how many ways we can choose one red, and two black divided by how many ways we can choose 3 marbles from 12. So: \[\frac{{7 \choose 1} * {5 \choose 2}}{12 \choose 3}\]

OpenStudy (anonymous):

so 12/3?

OpenStudy (anonymous):

No, 7 choose 1 = 7 5 choose 2 = 10 12 choose 3 = 220 So it is 70/220. Are you familiar with the choose function (and it's notation)?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

was that without replacement?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

do you know how to do d?

OpenStudy (anonymous):

For the other one it's just (Probability of choosing a red)*(Probability of choosing a black)*(Probability of choosing a black) So (7/12) * (5*/11) * (4/10)

OpenStudy (anonymous):

140/1320....can be reduced though right?

OpenStudy (anonymous):

14/132 = 7/66